OpenStudy (anonymous):

calculate the sum of 1/7 and 1/3

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\]

OpenStudy (anonymous):

to add fractions, they must have the same denominator to do this, you'll want to change the current denominators to a common multiple of them. multiplying them together will always give you a common multiple, but often, you can find a smaller one. in this case, the common multiple of 7 and 3 is 7x3 = 21 we want to add both fractions and we will make them /21. to do this we will multiply by what is really just ONE. \[\frac{ 1 }{ 7 } + \frac{ 1 }{ 3 } = \frac{ 1 }{ 7 }*\frac{ 3 }{ 3 } + \frac{ 1 }{ 3 } * \frac{ 7 }{ 7 } = \frac{ 3 }{ 21 } + \frac{ 7 }{ 21 } = \frac{ 3+7 }{ 21 } = \frac{ 10 }{ 21 }\]

OpenStudy (anonymous):